# Learning Latent Space Dynamics with Model-Form Uncertainties: A Stochastic Reduced-Order Modeling Approach

### v1

@@ 3.1.3 On Constructing the Global Projection Matrix
To capture the uncertainties raised by the training strategy for instance, one can consider $m$ combinations of snapshots corresponding to various initial conditions. In the OpInf framework where reduced-order operators are learned from data, it is necessary to identify these combinations such that the respective operators do not deviate significantly from each other—hence maintaining operator consistency across the anchor points. To achieve this, denote any of the matrix-valued parameters $\hat{\boldsymbol{c}}$, $[\hat{A}]$, $[\hat{H}]$ and $[\hat{P}]$, as $[O]$, and let $[D^{[O]}]$ be the pairwise Frobenius norm distance matrix for this operator, obtained over $N_{\text{comb}}$ combinations of parameters. Specifically, we define the entries of $[D^{[O]}]$ as
$D^{[O]}_{ij}=||[O^{(i)}]-[O^{(j)}]||_{F}\,,\quad 1\leq i,j\leq N_{\text{comb}}\,,$ (23)
@@
$\widetilde{D}^{[O]}_{ij}=\frac{D^{[O]}_{ij}}{\text{max}([D^{[O]}])-\text{min}([D^{[O]}])}\,,\quad 1\leq i,j\leq N_{\text{comb}}\,,$ (24)
@@
In order to ensure sufficient variability over the set of anchor points, we first perform Riemannian K-means clustering [55] on the set of all matrices $[\Phi^{(i)}]$ for $i=1,\dots,N_{\text{comb}}$ to identify $m\geq 3$ clusters. The clustering is performed on the subset of the Stiefel manifold $\mathbb{S}_{N,(r+q)}\subset\text{St}(N,(r+q))$. We utilize the iterative algorithm proposed in [22] to identify the optimal number of clusters $m$. For each cluster, we select as the anchor point $[\Phi^{(i)}]\in\mathcal{A}$ the basis matrix that minimizes the total pairwise distance to other matrices within the same cluster for all reduced-order operators, that is,
@@
When Riemannian K-means proves computationally intensive, due to numerous geodesic distance calculations and Riemannian exponential/logarithm operations, an alternative approach consists in performing K-means on the spectral embedding of the matrices $[\Phi^{(i)}]$ for $i=1,\dots,N_{\text{comb}}$, followed by standard K-means clustering on the resulting embedding. This approach offers a computationally efficient alternative while still capturing the underlying structure of the data on the Stiefel manifold.
@@ 3.2 Parameterization of the Probabilistic Model
$\boldsymbol{\alpha}=\text{arg}\underset{\boldsymbol{\alpha}\,\in\,\mathbb{R}^{m}_{>0}}{\text{min}}\boldsymbol{\alpha}^{\top}[H]\boldsymbol{\alpha}\,,$ (29)
where $[H]\in\mathbb{R}^{m\times m}$ is the symmetric positive-definite matrix, the entries of which depend on the anchor points through
$H_{ij}=\operatorname{tr}\left\{\log_{[\Phi^{*}]}^{\text{St}}\left([\Phi^{(i)}]\right)^{\top}\log_{[\Phi^{*}]}^{\text{St}}\left([\Phi^{(j)}]\right)\right\}\,,$ (30)
@@ 3.3 Selection of Reduced-Order Matrix Operators for Uncertainty Propagation
In order to propagate model-form uncertainties, Monte Carlo simulations are used in combination with the stochastic reduced-order model. Note that other non-intrusive stochastic solvers, such as stochastic collocation methods, can be used; see [20] for a review. The propagation task requires selecting appropriate reduced-order matrix operators for time integration in the reduced space for the $j$th sample $[\Phi(\theta_{j})]$ of $[\mathbf{\Phi}]$. Here, we employ a probabilistic interpretation leveraging the convex combination definition of the samples. Specifically, for each sample $[\Phi(\theta_{j})]$, we define an index $\mathcal{I}_{j}$, with $1\leq\mathcal{I}_{j}\leq m$, such that:
@@
This criterion associates $[\Phi(\theta_{j})]$ with the closest anchor point $[\Phi^{(\mathcal{I}_{j})}]$, operator-wise. Time integration in the reduced space is then carried out with the reduced-order matrix operators $\hat{\boldsymbol{c}}^{(\mathcal{I}_{j})}$, $[\hat{A}^{(\mathcal{I}_{j})}]$, $[\hat{H}^{(\mathcal{I}_{j})}]$, and $[\hat{P}^{(\mathcal{I}_{j})}]$ for the $j$th sample.
@@ 3.4 Summary of the Proposed Approach
The proposed stochastic reduced-order modeling approach begins with the generation of $N_{\text{comb}}$ snapshot matrices, followed by the completion of four subsequent tasks. First, projection matrices $[\Phi^{(i)}]$ and reduced-order matrix operators $\hat{\boldsymbol{c}}^{(i)}$, $[\hat{A}^{(i)}]$, $[\hat{H}^{(i)}]$, $[\hat{P}^{(i)}]$ are computed for $1\leq i\leq N_{\text{comb}}$ training scenarios. Second, anchor points are identified via clustering techniques, followed by the construction of a global projection matrix. Third, stochastic samples of the projection matrix are generated through a convex combination of anchor points on the tangent space, employing a Dirichlet distribution for the coefficients. Finally, appropriate reduced-order matrix operators are selected for uncertainty propagation based on a probabilistic interpretation of the stochastic samples. This approach is summarized in (pseudo-)Algorithm LABEL:alg:_SummaryAlgorithm.
@@ 4.1.1 Problem Description
and zero boundary conditions on the ends of $\Omega$. In this work, we consider $\operatorname{Re}=1,000$, which leads to sharp gradients in the solution and results in a problem where the Kolmogorov N-width decreases much slower than the commonly considered $\operatorname{Re}=100$ in the operator learning literature. This choice of Reynolds number is deliberate, as it creates a more challenging scenario for reduced-order modeling. The spatial domain $\Omega$ is discretized with $256$ Lagrange elements of degree 1. The step size used for time integration is $\Delta t=10^{-3}$ and a total of $k=2001$ snapshots are collected for each value of $\mu$ for training the operators. We rely on implicit backward Euler scheme in conjunction with a Newton-Raphson solver for time integration. We collected snapshots for different initial condition parameter $\mu=[0.4,0.5,\dots,1.2]$ with an interval of 0.1 and subsequently constructed multiple training datasets based on combinations of parameter $\mu$.
@@
The full-order solution of the Burgers’ equation (32) at various time instances for different initial condition parameter $\mu$ is shown in Fig. 2.
@@ 4.1.2 Parametric Analysis on Reduced-Order Model
It is observed that MPOD-OpInfPoly have a better relative state error compared to the POD-OpInf model at lower values of $r$. This is attributed to the polynomial representation of the operators, which allows for a more accurate representation of the dynamics. However, as the number of reduced basis vectors $r$ increases, the relative state error of the POD-OpInf model decreases and eventually converges to the MPOD-OpInfPoly model. This suggests that the polynomial representation of the operators is not necessary when a sufficient number of reduced basis vectors are used. Outside the range of training data ($t\in[2,8]$), the relative state error of the MPOD-OpInfPoly model is higher than the MPOD-OpInf model at lower values of $r$, which could be attributed to the overfitting of the polynomial representation. These results are consistent with the observations reported in [18]. Note that better results are obtained in testing for the POD-OpInf procedure. A possible explanation is that the testing region (from $t=2$ to $t=8$) mainly comprises diffusive patterns which are easily captured by the dominant POD modes, whereas the training region contains shocks or sharper gradients that might require a larger number of POD modes to resolve.
@@ 4.1.3 Uncertainty Quantification Results
Figure 5: Graph of the error function $r\mapsto\epsilon(r)$ for different combinations of parameter $\mu$. The vertical axis represents the error function $\epsilon(r)$ as defined in Eq. (34). The horizontal axis shows the rank $r$ of the projection matrices $[\Phi^{(i)}]$, where $1\leq i\leq N_{\text{comb}}$, and $N_{\text{comb}}$ is the number of parameter combinations.
@@ 4.2.2 Uncertainty Quantification Results
Figure 11: Graph of the error function $r\mapsto\epsilon(r)$ for different combinations of parameter $\mu$ for the two-dimensional transient flow past circular cylinder problem. The horizontal axis shows the rank $r$ of the projection matrices $[\Phi^{(i)}]$, where $1\leq i\leq N_{\text{comb}}$, and $N_{\text{comb}}$ is the number of parameter combinations.

### v2

@@ 3.1.3 On Constructing the Global Projection Matrix
To capture the uncertainties raised by the training strategy for instance, one can consider $m$ combinations of snapshots corresponding to various initial conditions. In the OpInf framework where reduced-order operators are learned from data, it is necessary to identify these combinations such that the respective operators do not deviate significantly from each other—hence maintaining operator consistency across the anchor points. To achieve this, denote any of the matrix-valued parameters $\hat{\boldsymbol{c}}$, $[\hat{A}]$, $[\hat{H}]$ and $[\hat{P}]$, as $[O]$, and let $[D^{[O]}]$ be the pairwise Frobenius norm distance matrix for this operator, obtained over $n_{c}$ combinations of parameters. Specifically, we define the entries of $[D^{[O]}]$ as
$D^{[O]}_{ij}=||[O^{(i)}]-[O^{(j)}]||_{F}\,,\quad 1\leq i,j\leq n_{c}\,,$ (23)
@@
$\widetilde{D}^{[O]}_{ij}=\frac{D^{[O]}_{ij}}{\text{max}([D^{[O]}])-\text{min}([D^{[O]}])}\,,\quad 1\leq i,j\leq n_{c}\,,$ (24)
@@
In order to ensure sufficient variability over the set of anchor points, we first perform Riemannian K-means clustering [62] on the set of all matrices $[\Phi^{(i)}]$ for $i=1,\dots,n_{c}$ to identify $m\geq 3$ clusters. The clustering is performed on the subset of the Stiefel manifold $\mathbb{S}_{N,(r+q)}\subset\text{St}(N,(r+q))$. We utilize the iterative algorithm proposed in [24] to identify the optimal number of clusters $n_{c}$. For each cluster, we select as the anchor point $[\Phi^{(i)}]\in\mathcal{A}$ the basis matrix that minimizes the total pairwise distance to other matrices within the same cluster for all reduced-order operators, that is,
@@
When Riemannian K-means proves computationally intensive, due to numerous geodesic distance calculations and Riemannian exponential/logarithm operations, an alternative approach consists in performing K-means on the spectral embedding of the matrices $[\Phi^{(i)}]$ for $i=1,\dots,n_{c}$, followed by standard K-means clustering on the resulting embedding. This approach offers a computationally efficient alternative while still capturing the underlying structure of the data on the Stiefel manifold.
@@ 3.2 Parameterization of the Probabilistic Model
$\boldsymbol{\alpha}=\text{arg}\underset{\boldsymbol{\alpha}\,\in\,\mathbb{R}^{m}_{>0}}{\text{min}}\boldsymbol{\alpha}^{\top}[Q]\boldsymbol{\alpha}\,,$ (29)
where $[Q]\in\mathbb{R}^{m\times m}$ is the symmetric positive-definite matrix, the entries of which depend on the anchor points through
$Q_{ij}=\operatorname{tr}\left\{\log_{[\Phi^{*}]}^{\text{St}}\left([\Phi^{(i)}]\right)^{\top}\log_{[\Phi^{*}]}^{\text{St}}\left([\Phi^{(j)}]\right)\right\}\,,$ (30)
@@ 3.3 Selection of Reduced-Order Matrix Operators for Uncertainty Propagation
In order to propagate model-form uncertainties, Monte Carlo simulations are used in combination with the stochastic reduced-order model. Note that other non-intrusive stochastic solvers, such as stochastic collocation methods, can be used; see [22] for a review. Observe that while all reference “models” (encoded through the projection matrices $[\Phi^{(i)}]$, $1\leq i\leq n_{c}$), obtained by running over all possible combinations of training parameters, are associated with their corresponding reduced-order matrix operators, the new samples for the projection matrix (i.e., the $j$th sample $[\Phi(\theta_{j})]$ of $[\mathbf{\Phi}]$) are not associated with any such matrices. Hence, the approach requires the identification of appropriate reduced-order matrix operators for a given sample $[\Phi(\theta_{j})]$, so as to minimize bias during time integration in the reduced space. We proceed to introduce a criterion for the selection of suitable reduced-order matrix operators.
Here, we employ a probabilistic interpretation leveraging the convex combination definition of the samples. Specifically, for each sample $[\Phi(\theta_{j})]$, we define an index $\mathcal{I}_{j}$, with $1\leq\mathcal{I}_{j}\leq m$, such that:
@@
This criterion associates $[\Phi(\theta_{j})]$ with the closest anchor point $[\Phi^{(\mathcal{I}_{j})}]$, operator-wise. Time integration in the reduced space is then carried out with the selected reduced-order matrix operators $\hat{\boldsymbol{c}}^{(\mathcal{I}_{j})}$, $[\hat{A}^{(\mathcal{I}_{j})}]$, $[\hat{H}^{(\mathcal{I}_{j})}]$, and $[\hat{P}^{(\mathcal{I}_{j})}]$ for the $j$th sample.
@@ 3.4 Summary of the Proposed Approach
The proposed stochastic reduced-order modeling approach begins with the generation of $n_{c}$ snapshot matrices, followed by the completion of four subsequent tasks. First, projection matrices $[\Phi^{(i)}]$ and reduced-order matrix operators $\hat{\boldsymbol{c}}^{(i)}$, $[\hat{A}^{(i)}]$, $[\hat{H}^{(i)}]$, $[\hat{P}^{(i)}]$ are computed for $1\leq i\leq n_{c}$ training scenarios. Second, $m$ anchor points are identified via clustering techniques, followed by the construction of a global projection matrix. Third, stochastic samples of the projection matrix are generated through a convex combination of anchor points on the tangent space, employing a Dirichlet distribution for the coefficients. Finally, appropriate reduced-order matrix operators are selected for uncertainty propagation based on a probabilistic interpretation of the stochastic samples. This approach is summarized in (pseudo-)Algorithm LABEL:alg:_SummaryAlgorithm.
@@ 4.1.1 Problem Description
and zero boundary conditions on the ends of $\Omega$. In this work, we consider $\operatorname{Re}=1,000$, which leads to sharp gradients in the solution and results in a problem where the Kolmogorov N-width decreases much slower than the commonly considered $\operatorname{Re}=100$ in the operator learning literature. This choice of Reynolds number is deliberate, as it creates a more challenging scenario for reduced-order modeling.
To collect snapshots for training the operators, we consider the range $t\in[0,2]$ for training data, and discretize the spatial domain $\Omega$ with $256$ $\text{P}_{1}$ Lagrange elements. The step size used for time integration is $\Delta t=10^{-3}$ and a total of $k=2001$ snapshots are collected for each value of $\mu$ for training the operators. We rely on implicit backward Euler scheme in conjunction with a Newton-Raphson solver for time integration. We collected snapshots for different initial condition parameter $\mu=[0.4,0.5,\dots,1.2]$ with an interval of 0.1 and subsequently constructed multiple training datasets based on combinations of parameter $\mu$.
@@
The full-order solution of the Burgers’ equation (32) at various time instances for different initial condition parameter $\mu$ is shown in Fig. 2. The testing window employed for this problem is $t\in[2,8]$.
@@ 4.1.2 Parametric Analysis on Reduced-Order Model
It is observed that MPOD-OpInfPoly has a better relative state error compared to the POD-OpInf model at lower values of $r$. This is attributed to the polynomial representation of the operators, which allows for a more accurate representation of the dynamics. However, as the number of reduced basis vectors $r$ increases, the relative state error of the POD-OpInf model decreases and eventually converges to the MPOD-OpInfPoly model. This suggests that the polynomial representation of the operators is not necessary when a sufficient number of reduced basis vectors are used. Outside the range of training data ($t\in[2,8]$), the relative state error of the MPOD-OpInfPoly model is higher than the MPOD-OpInf model at lower values of $r$, which could be attributed to the overfitting of the polynomial representation. These results are consistent with the observations reported in [20]. Note that better results are obtained in testing for the POD-OpInf procedure. A possible explanation is that the testing region (from $t=2$ to $t=8$) mainly comprises diffusive patterns which are easily captured by the dominant POD modes, whereas the training region contains shocks or sharper gradients that might require a larger number of POD modes to resolve.
@@ 4.1.3 Uncertainty Quantification Results
Figure 5: Graph of the error function $r\mapsto\epsilon(r)$ for different combinations of parameter $\mu$. The vertical axis represents the error function $\epsilon(r)$ as defined in Eq. (34). The horizontal axis shows the rank $r$ of the projection matrices $[\Phi^{(i)}]$, where $1\leq i\leq n_{c}$, and $n_{c}$ is the number of parameter combinations. Distinct colors indicate different combination groups ${}^{9}C_{k}$ (where $3\leq k\leq 9$ and ${}^{n}C_{k}$ denotes $k$-combinations from a set of $n$ elements). The number of combinations within each group is denoted by $n$ in the figure.
@@ 4.2.2 Uncertainty Quantification Results
Figure 11: Graph of the error function $r\mapsto\epsilon(r)$ for different combinations of parameter $\mu$ for the two-dimensional transient flow past circular cylinder problem. The horizontal axis shows the rank $r$ of the projection matrices $[\Phi^{(i)}]$, where $1\leq i\leq n_{c}$, and $n_{c}$ is the number of parameter combinations. Distinct colors indicate different combination groups ${}^{9}C_{k}$ (where $3\leq k\leq 9$ and ${}^{n}C_{k}$ denotes $k$-combinations from a set of $n$ elements). The number of combinations within each group is denoted by $n$ in the figure